OpenStudy (anonymous):

Can someone help me figure this out? Select one of the factors of x^2 – 17x + 72 I will leave the possible answers here as reference: a)x - 8 b)x - 2 c)x + 9 d)x + 12 I am taking Algebra 2 and am currently pulling a high "B". I want to bring that up to an "A". Can anyone explain the steps I need to figure this out? :D

jhonyy9 (jhonyy9):

(x-8)(x-9)

OpenStudy (anonymous):

How did you figure that out?

OpenStudy (jlastino):

You should two numbers whose product is 72 And whose sum is 17

OpenStudy (jlastino):

*find

OpenStudy (anonymous):

x^2-17x+72=0 x^2-8x-9x+72=0 x(x-8)-9(x-8)=0 (x-8)(x-9)=0

OpenStudy (anonymous):

Alright, that makes sense. Thanks! I just got the same answer! Thank you so much!

OpenStudy (anonymous):

Ok, so I have this problem now which is 4x^2 - 8x - 140. A) x + 5 B) 8 C) x - 4 D) x - 2 I tried getting the answer by finding two numbers that get the product of -140 and the sum of -8. I ended up with (x + 5)(x - 13) so that would give me -8 but not a product of -140. I am hopelessly confused, can anyone clear this up for me?

OpenStudy (jlastino):

You can simplify that by taking out the common factor "4"

OpenStudy (anonymous):

So, it would end up being 4(x + y)(x + y)

OpenStudy (anonymous):

y being a placeholder

OpenStudy (anonymous):

Divide through out by 4 you get x^2-2x-35=0 x^2-7x+5x-35=0 X(x-7)-5(x-7)=0 (x-7)(x-5)=0

OpenStudy (jlastino):

yeah but first find what the quad equation would look like

OpenStudy (anonymous):

*x+5

OpenStudy (anonymous):

Factoring equations sucks.. :(

OpenStudy (jlastino):

scasm already did the work for you though

OpenStudy (anonymous):

I see that, I was trying to figure it out so I didn't have to keep asking for help. :P

OpenStudy (jlastino):

you'll eventually get used to it :)) its really frustrating at first you know

OpenStudy (anonymous):

Alright, I think this will be the last question I need to understand it. 2x^2 + 16x - 18 I factored out the common factor which is 2. so now I have x^2 + 8x - 9 I imagine the quad would be x^2 + 4x + 4x - 9. what can I do from here or what did I do wrong?

OpenStudy (anonymous):

You dont get 4 and 4. Since the last number is 9, u have this with some permutations and combinations get two numbers such that their product is 9 and their sum is 8. Here 3*3*1 are factors of 9 .So its 9 and 1 equation will be x^2+9x-x-9=0 Here the product should be -9 as indicated by the equation, so the factors are 9 and -1 which also complies with their sum 9-1=8 which is right in the middle of the eqn